You want to buy a house that costs $200,000 and have saved up enough for the 10% down payment. You will be borrowing the rest from the bank at an annual rate of 9% compounded s.a. through a 25 year mortgage.

  • How much will your monthly payments be?
  • How much of the first monthly payment will go towards principal?
  • What will be the total cost of your house?
  • How much remains owing at the end of the 3 years, and what percentage of your first 36 payments went to principal?
Answer #1

Loan Amount = (1 - 0.10)(200,000) = $180,000

Calculating Monthly Payment on Loan,

Using TVM Calculation,

PMT = [PV = 180,000, FV = 0, N = 300, I = 0.09/12]

PMT = $1,510.55

Principal Amount in First Payment = 1,510.55 - 180,000(0.09)/12

Principal Amount in First Payment = $160.55

Total Cost = (1,510.55)(300) + 20,000 = $473,165